OpenStudy (anonymous):

Helppp Pleeaassee!!! Estimate the value of the logarithm to the nearest tenth log9 61 a. 6.3 b. 0.5 c. -1.9 d. 1.9

Nnesha (nnesha):

\[\huge\rm log_a b = \frac{ \log b }{ \log a }\] change of base formula apply that

OpenStudy (anonymous):

ok.. how?

Nnesha (nnesha):

like this \[\huge\rm log_a b = \frac{ \log b }{ \log a }\]